Question:

The mean score on an accounting test is 60, with a standard deviation of 8. Between which two scores must this mean lie to represent 8/9 of the data set?

Find first the value of k.


Use Chebychev’s Inequality to solve this.



1
Expert's answer
2021-12-23T04:54:12-0500

Mean = 60

Standard deviation = 8

The value of k as per Chebychev’s Inequality rule:


"1-\frac{1}{k^2}=\frac{8}{9}"


"1-\frac{8}{9}=\frac{1}{k^2}"


"\frac{1}{9}=\frac{1}{k^2}"


"k^2=9"


"k=3"


Lower score limit:

"60-(k)(8)=60-(3\times8)=36"


Upper score limit:

"60+(k)(8)=60+(3\times8)=84"


The mean score on an accounting test is 60, with an standard deviation of 8 must lie between 36 and 84 to represent 8/9 of the data set.